MATH 230 SYLLABUS (REVISED) Spring 1994

Section: 230.3; Monday, Wednesday, Thursday, Friday, 11:15 AM to 12:05 PM; 273 Willard. Instructor: Stephen G. Simpson; 333 McAllister; 863–0775. Office Hours: Mondays 3:15–4:30 PM, Thursdays 2:45-4:30 PM. NOTE: Stephen G. Simpson is not to be confused with Todd A. Simpson, the instructor of another section of Math 230. Course Description: MATH 230. CALCULUS AND VECTOR ANALYSIS (4 credits). Three-dimensional analytic geometry; vectors in space; partial differentiation; double and triple integrals; integral vector calculus. Prerequisite: MATH 141. Textbook: C. H. Edwards, Jr. and D. E. Penney, Calculus and Analytic Geometry, Third Edition, Prentice– Hall, 1990.

January 10–14. §§13.3–13.5. Vectors in the plane; position, velocity, and acceleration vectors; projectiles; polar coordinates; radial and transverse components of velocity and acceleration; Kepler’s laws. QUIZ #1, Monday, January 17. January 17–21. §§14.1–14.4. Vectors in space; vector products; lines and planes in space; curves and moving points in space. January 24–28. §§14.5, 14.7. Curvature; normal component of acceleration; cylindrical and spherical coordinates. QUIZ #2, Monday, January 31. January 31 – February 4. §§15.1–15.4, 15.5, 15.10. Partial derivatives; maxima and minima for functions of several variables; second derivative test. FIRST MIDTERM EXAM, Wednesday, February 9, 6:30–7:45 PM February 7–11. §§15.6, 15.7. Differentials, chain rule; gradients; directional derivatives. QUIZ #3, Monday, February 14. February 14–18. §§15.8, 15.9. More on maxima and minima; Lagrange multipliers. February 21–25. §§16.1, 16.2. Double integrals; double integrals over general regions. QUIZ #4, Monday, February 28. February 28 – March 4. §§16.3, 16.4. Double integrals; area and volume; polar coordinates; change of variables. March 7–11. SPRING BREAK. March 14–18. §§16.5, 16.8, 16.9. Surface area; change of variables.

SECOND MIDTERM EXAM, Monday, March 21, 6:30–7:45 PM March 21–25. §§16.6, 16.7, 16.9. Triple integrals. Cylindrical and spherical coordinates; change of variables. March 28–April 1. §§17.1, 17.2, 17.3. Vector fields; line integrals; potential functions. QUIZ #5, Monday, April 4. April 4–8. §§17.3, 17.4. Path-independent line integrals; Green’s theorem. April 11–15. §§17.5, 17.6. Surface integrals; orientable surfaces; the Divergence theorem. QUIZ #6, Monday, April 18. April 18–22. §§17.6, 17.7. Stokes’ theorem. Examples. April 25–29. Catch up and review. FINAL EXAM (date and time to be announced).

Grades:  Quizzes, homework  100
Midterm Exams  2 × 100 = 200
Final Exam  150
Total  450